OpenStudy (anonymous):

Jill took five courses this semester, each for four credit hours. She received a B (3 points) a B+ (3.5 points), an A (4 points), and a B+ (3.5 points) in four of the courses. If her GPA is 3.2, what was her grade in the fifth course?

OpenStudy (anonymous):

she has a b

OpenStudy (anonymous):

14 + x = 5(3.2) x = 16 - 14 x = 2.0 for the 5th course

OpenStudy (anonymous):

yaya!

OpenStudy (anonymous):

3+3.5+ 4++3.5+x= 3.2 ------------------- 5 * both sides by 5, + the 4 grades

OpenStudy (anonymous):

did i get it right you think?

OpenStudy (anonymous):

? yes

OpenStudy (anonymous):

VERYYYYYYYYYYY CORRECT

OpenStudy (anonymous):

jayyy thanks :)